OpenStudy (anonymous):

Does anyone know how to compute 28P7?? 0.o

terenzreignz (terenzreignz):

In fact, there is a formula for nPr, it goes : \[\Large _nP_r = \frac{n!}{(n-r)!}\]

terenzreignz (terenzreignz):

But the layman's definition would probably more along the lines of... \[\Large _{28}P_7 = 28\cdot 27 \cdot ... [\text{until you have a total of 7 factors}]\]

OpenStudy (anonymous):

so it would be 28!/(28-7)! ???

terenzreignz (terenzreignz):

Yes, on the whole... yes :D

OpenStudy (anonymous):

Which is 28!/21! ??? Right?

terenzreignz (terenzreignz):

Yep.

OpenStudy (anonymous):

Ohhhhh ok I get it so the final answer is going to be really long lol

terenzreignz (terenzreignz):

yup... don't be surprised if you get ridiculously large numbers when dealing with these permutation types...

OpenStudy (anonymous):

Ok so bear with me on this one. It would be (28 x 27 x 26 x 25 x 24 x 23 x 22 x 21!) (21!)

OpenStudy (anonymous):

28 x 27 x 26 x 25 x 24 x 23 x 22

terenzreignz (terenzreignz):

You missed a slash / You seriously don't want that number to be bigger than it should be, it'd blow up your calculator... but, it's right :D

OpenStudy (anonymous):

WHICH ISSSSSSSSSS 5967561600!!!!

terenzreignz (terenzreignz):

Did you notice that that's exactly how I described it? 28 x 27 x ... until you have 7 factors? 28 x 27 x 26 x 25 x 24 x 23 x 22 --> 7 factors :P

OpenStudy (anonymous):

Yeah your way was way easier lol thanks bro

terenzreignz (terenzreignz):

No problem ^_^ The name's TJ by the way :D

OpenStudy (anonymous):

Thanks TJ